### v1

# Heterogeneous Mixed Traffic Control and Coordination

Abstract
Urban intersections, filled with a diverse mix of vehicles from small cars to large semi-trailers, present a persistent challenge for traffic control and management. This reality drives our investigation into how robot vehicles (RVs) can transform such heterogeneous traffic flow, particularly at unsignalized intersections where traditional control methods often falter during power failures and emergencies. Using reinforcement learning (RL) and real-world traffic data, we study heterogeneous mixed traffic across complex intersections under gradual automation by varying RV penetration from 10% to 90%. The results are compelling: average waiting times decrease by up to 86% and 91% compared to signalized and unsignalized intersections, respectively. Additionally, we uncover a “rarity advantage,” where less frequent vehicles, such as trucks, benefit the most from RV coordination (by up to 87%). RVs’ presence also leads to lower CO2 emissions and fuel consumption compared to managing traffic via traffic lights. Moreover, space headways decrease across all vehicle types as RV rate increases, indicating better road space utilization. These results highlight the potential of RVs to manage heterogeneous mixed traffic in complex environments with optimized traffic efficiency and minimized environmental impact.
I INTRODUCTION
Urban roadways are dominated by heterogeneous traffic, characterized by a mix of vehicles such as passenger cars, trucks, vans, and more. The complexities arising from the varying sizes, speeds, and maneuverability of these vehicle types significantly affect overall road performance [1]. For example, larger vehicles like trucks require more space and tend to cause delays, while smaller cars navigate more easily [2, 3]. These differences often disrupt traffic flow, as slower vehicles interfere with the movement of faster ones, increasing instability and complicating traffic management [4]. As a result, traffic congestion intensifies, causing extended commute times and significant economic impacts. In the U.S. alone, drivers lose an average of 100 hours annually to traffic congestion, resulting in economic losses of \$1,400 per person [5].
Intersections play a crucial role in urban traffic systems where vehicles converge and disperse. Traditionally, traffic lights are used as to manage these complex interactions. However, intersections can turn into congestion points, especially when such control methods fail or are absent. Natural disasters, large-scale power failures, or system malfunctions can disable traffic signals, converting typically controlled intersections into unregulated crossings. Additionally, during temporary road configuration due to construction, intersections may lack signalization. These various situations pose significant traffic management challenges for urban areas, forcing them to navigate complex traffic conditions without traffic lights for extended periods [6, 7].
The emergence of robot vehicles (RVs) with varying degrees of self-driving capabilities has the potential to transform urban traffic management. These vehicles can be leveraged to regulate nearby human-driven vehicles (HVs) so that the mixed traffic, consisting of both RVs and HVs, can be optimized for various objectives such as improving efficiency or reducing emissions [8, 9, 10, 11, 12, 13, 14]. When selecting control strategies for mixed traffic, reinforcement learning (RL) has demonstrated significant potential due to its model-free characteristics [13, 15, 16, 17]. RL is particularly well-suited for managing multi-agent systems at complex and dynamic environments, e.g., traffic at unsignalized intersections, because of its flexibility to adapt and learn effective strategies [18]. Unlike traditional rule-based systems, RL can potentially discover novel solutions that might not be apparent to human traffic engineers [19]. This combination of RVs and RL presents a promising approach for addressing the challenges of mixed traffic control and coordination.
Despite the aforementioned advancements, a significant gap exists between current research of mixed traffic and the complexities of real-world traffic: nearly all existing studies assume homogeneous traffic. Real-world traffic, on the other hand, features a variety of vehicle types, i.e., heterogeneous traffic, equipped with a wide range of kinematic and dynamic parameters. These factors are essential for effective traffic control and management [1, 2, 3, 4]. The limitations of homogeneous traffic become more pronounced when factoring in the complexity of test environments and varying RV penetration rates, raising several important research questions.
  • Is it feasible to implement heterogeneous mixed traffic control in complex environments?
  • What are the performance implications for entire traffic and different vehicle types in terms of efficiency and environmental impact?
  • How do varying RV penetration rates affect performance outcomes?
Addressing these questions will significantly enhance the applicability of mixed traffic control and coordination, particularly in complex environments like unsignalized intersections where traditional control methods may be absent.
In this study, we aim to fill the critical gap by demonstrating the feasibility of heterogeneous mixed traffic control and coordination in complex environments. We also conduct comprehensive experiments to evaluate efficiency and environmental impact of the heterogeneous traffic. Specifically, our contributions are the following.
  • Our approach incorporates multiple vehicle types, including passenger cars, pickups, vans, semi-trailers, and trucks, extracted from real-world traffic data.
  • Our method is implemented and tested across multiple complex, real-world intersection layouts.
  • Our experiments are grounded in vehicle type distributions that mirror actual urban traffic composition.
  • Our detailed analysis examines various RV penetration rates to understand the impact of gradual automation in mixed traffic environments.
To the best of our knowledge, this is the first study investigating the RL-based control and coordination of heterogeneous mixed traffic in complex environments. The findings and insights extend the frontier of emerging mixed traffic research.
II RELATED WORK
The diversity in vehicle characteristics within heterogeneous traffic poses numerous challenges for traffic management. Studies show that the composition of traffic significantly impacts overall road performance, with larger and slower vehicles often exacerbating congestion and causing delays [1, 20]. Sun et al. [4] demonstrate that considering traffic heterogeneity, particularly leader-follower compositions and driving styles, is crucial for accurately modeling traffic dynamics, as neglecting these factors can lead to substantial estimation errors.
At intersections, these challenges are more pronounced. Mohan and Chandra [3] reveal that larger vehicles like trucks and buses require longer gaps to navigate intersections safely, which reduces overall intersection capacity and increases delays. Furthermore, safety evaluations and conflicting vehicle speeds indicate that interactions between diverse vehicle types greatly affect the likelihood of conflicts [21]. Suriyarachchi et al. [22] propose a game-theoretic framework to manage heterogeneous traffic at intersections. Their approach optimizes the sequence of vehicle entries at unsignalized intersections to reduce fuel consumption. However, game-theoretic methods may face scalability issues in more unpredictable traffic scenarios [23].
Several studies explore the management of traffic at unsignalized intersections, particularly for connected and autonomous vehicles (CAVs). Early work by Dresner and Stone [24] introduces a multi-agent intersection control mechanism where CAVs reserve space-time slots in a first-come, first-serve (FCFS) manner. Jin et al. [25] expand on this pioneer work by applying FCFS scheduling to CAV platoons, while Chen et al. [26] propose a controllable gap strategy that adjusts the time gap between vehicles based on their speed and conflict relationships to prevent collisions. Researchers also examine decentralized approaches, such as the consensus-based control method for managing the trajectories of CAVs by Mirheli et al. [27] and Malikopoulos et al.’s work [28] for optimizing the energy consumption. These methods typically assume a fully autonomous environment.
Recently, studies show the effectiveness of RL in managing and optimizing mixed traffic across various scenarios. Wang et al. [17, 29] demonstrate the effectiveness of decentralized multi-agent RL approach for managing mixed traffic at complex, unsignalized intersections, showing significant reductions in waiting time with high RV penetration. Villarreal et al. [10] explore the use of image observations instead of precise traffic information for mixed traffic control via RL. Poudel et al. [30, 31] propose an RL framework that incorporates real-world driving profiles into mixed traffic control and enhances traffic safety, stability, and efficiency. Yan and Wu [32] introduce a model-free RL method for controlling mixed traffic at unsignalized intersections using multi-agent policy decomposition. Peng et al. [33] apply RL to improve mixed traffic efficiency at unsignalized intersections via CAVs. Similarly, Shi et al. [34] present how RL-based method can effectively reduce collisions and improve traffic flow at urban unsignalized intersections compared to traditional control methods.
Despite significant progress, most existing studies face two major limitations: 1) homogeneous traffic is assumed, and 2) the test environment is simplified (either by limiting the number of conflicting vehicles or using idealized intersection topologies). This lack of real-world complexity hinders the applicability of the proposed techniques. We aim to fill the gap by demonstrating the effectiveness of RL for managing heterogeneous mixed traffic in complex environments.
III METHODOLOGY
III-A RL-based Heterogeneous Traffic Control
We formulate the problem of managing heterogeneous mixed traffic at unsignalized intersections as a Partially Observable Markov Decision Process (POMDP), defined as the collection $(\mathcal{X},\mathcal{U},\mathcal{P},\mathcal{R},\mathcal{Y},\mathcal{V},\mathcal{P}_{0},\gamma)$. Here, $\mathcal{X}$ represents the set of all possible states; $\mathcal{U}$ denotes the set of available actions; $\mathcal{P}({x}^{\prime}|{x},u)$ describes the probability of transitioning to a new state ${x}^{\prime}$ given the present state ${x}$ and action $u$; and $\mathcal{R}({x},u)$ defines the reward function that provides feedback based on the state-action pair. The observation space $\mathcal{Y}$ captures the partial view of the environment each RV receives, with $\mathcal{V}(y|{x})$ representing the probability of receiving an observation $y$ from the state ${x}$. $\mathcal{P}_{0}$ is the distribution of initial states, and $\gamma$ represents the factor for discounting future rewards. At each step $t$, an RV chooses an action $u_{t}\in\mathcal{U}$ using the policy $\pi_{\theta}(u_{t}|{x}_{t})$. The environment transitions to the next state ${x}_{t+1}$, and the RV receives a reward $r_{t}$. The goal is to maximize the total discounted reward: $\mathcal{R}_{t}=\sum_{i=t}^{T}\gamma^{i-t}r_{i}$ [17].
RVs decide between two possible actions: Stop and Go. Stop halts the RV before entering the intersection, while Go allows it to proceed. The observation $y_{t}$ of an RV at time $t$ encompasses traffic conditions at the intersection:
$y_{t}=\bigoplus_{k\in K}\langle l_{k},w_{k}\rangle\bigoplus_{k\in K}o_{k},$ (1)
where $K$ is the set of traffic directions, $l_{k}$ is the queue length in direction $k$, $w_{k}$ is the average waiting time in direction $k$, and $o_{k}$ shows the intersection occupancy status.
We design the reward function to encourage traffic efficiency (reduced average waiting times) and discourage vehicle movement conflicts. At time step $t$, the reward $r_{t}$ is defined as $r_{t}=\alpha h_{t}+c_{t}$, where $h_{t}$ is the local reward, $\alpha$ is the weight for local rewards, and $c_{t}$ is the conflict penalty. $h_{t}$ is defined as:
$h_{t}=\begin{cases}-w_{k},&\text{if Stop};\\ w_{k},&\text{if Go},\end{cases}$
and $c_{t}$ is defined as:
$c_{t}=\begin{cases}-1,&\text{if conflict occurs};\\ 0,&\text{otherwise}.\end{cases}$
III-B Heterogeneous Traffic and Complex Intersections
To construct heterogeneous traffic, we analyze the real-world traffic dataset I-24 MOTION [35] and extract the vehicle types and their ratios within: passenger cars (70%), pickups (3%), vans (15%), semi-trailers (11%), and trucks (1%) [31]. The kinematic and dynamic parameters of these vehicle types are presented in Table I [35, 36]. Our study scenarios are four real-world intersections shown in Fig. 1, with the number of incoming lanes 21, 19, 18, and 16 for intersection I, II, III, and IV, respectively.
TABLE I: The kinematic and dynamic parameters of each vehicle type.
Parameter Passenger Car Pickup Van Semi-trailer Truck
Length (m) 5 5.8 5.5 16.5 7.1
Width (m) 1.8 2 2 2.55 2.4
Height (m) 1.5 1.9 2.1 4 2.4
Mass (kg) 1500 2500 3000 15000 12000
Min Gap (m) 2.5 2.5 2.5 2.5 2.5
Accel. (m/s2) 2.6 2.6 2.6 1.0 1.3
Decel. (m/s2) 4.5 4.5 4.5 4 4
Max Speed (m/s) 55.56 33.33 27.78 36.11 36.11
Figure 1: Real-world and complex intersections used in our study.
III-C Evaluation Metrics
We provide comprehensive analysis of heterogeneous mixed traffic in terms of traffic efficiency, environmental impact, and vehicle distribution using the following metrics.
III-C1 Average Waiting Time ($W_{\text{avg}}$)
Average waiting time is defined as the mean time that vehicles spend waiting at the intersection [17]: $W_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}W_{i}$, where $N$ is the total number of vehicles and $W_{i}$ is the waiting time of vehicle $i$. We use this metric to assess the overall traffic efficiency. Lower values indicate smoother traffic flow and reduced congestion.
III-C2 CO2 Emission ($E_{\text{CO}_{2}}$)
[…]
III-C3 Fuel Consumption ($F_{\text{avg}}$)
Fuel consumption is the average amount of fuel consumed by vehicles within the intersection over a specified time period [37]: $F_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}F_{i}$. $N$ is the total number of vehicles and $F_{i}$ is the fuel consumed by vehicle $i$. This metric provides insights into the energy efficiency of the traffic system.
III-C4 Space Headway ($H_{\text{space}}$)
Space headway is measured as the average distance between the front bumper of a vehicle and the front bumper of the vehicle immediately ahead [38]: $H_{\text{space}}=\frac{1}{M}\sum_{j=1}^{M}(X_{j+1}-X_{j})$, where $M$ is the total number of vehicle pairs, $X_{j}$ is the position of the vehicle $j$, and $X_{j+1}$ is the position of the vehicle immediately ahead of the vehicle $j$. Space headway represents how efficiently vehicles utilize available road space and maintain safe distances.
IV EXPERIMENTS
We begin by introducing our experiment setup, then present the results on overall traffic efficiency. This is followed by a detailed discussion of both efficiency and environmental impact within heterogeneous mixed traffic.
IV-A Experiment Setup
We model RVs as passenger cars to streamline the complexity of the simulations while still capturing the essence of mixed traffic dynamics. Passenger cars represent the most common vehicle type on urban roads, making them a practical choice for evaluating RV control and coordination. To evaluate the performance of heterogeneous mixed traffic, we compare two baselines: (1) HV-Signalized: all HVs with traffic signals, representing the traditional intersection management method; and (2) HV-Unsignalized: all HVs without traffic signals, representing the unsignalized scenarios. The selected metrics—average waiting time, CO2 emissions, fuel consumption, and space headway—enable us to holistically assess traffic efficiency and environmental impact.
For training the RL policy, we use Rainbow DQN [39] with 1,000 iterations, leveraging an Intel i9-13900KF processor and NVIDIA GeForce RTX 4090 graphics card. The policy’s architecture is a three-layer neural network with 512 units per hidden layer. The learning process uses a rate of 0.0005 and a 0.99 discount factor. RV penetration rates vary from 10% to 90%, incrementing by 10%. We perform 10 evaluations for each RL policy, averaging the outcomes over a 1000-second timeframe. The vehicle type ratios are adjusted based on the RV rates as shown in Table II.
TABLE II: Vehicle Type Ratios for Different RV Penetration Rates. The initial distribution is extracted from the I-24 MOTION dataset [35], with RVs introduced as passenger cars. As RV rate increases, only the passenger car (non-RV) ratio decreases until 70%, after which other vehicle type ratios are proportionally reduced to maintain the total at 100%.
RV (%) Passenger Car (non-RV) (%) Pickup (%) Van (%) Semi-trailer (%) Truck (%)
10 60 3 15 11 1
20 50 3 15 11 1
30 40 3 15 11 1
40 30 3 15 11 1
50 20 3 15 11 1
60 10 3 15 11 1
70 0 3 15 11 1
80 0 2 10 7.3 0.7
90 0 1 5 3.7 0.3
IV-B Overall Traffic Efficiency
The traffic demands for the four unsignalized intersections are extracted from real-world data [17]. Specifically, intersection I has 1,157 vehicles per hour (v/h), intersection II has 1,089 v/h, intersection III sees 928 v/h, and intersection IV experiences the lowest demand at 789 v/h.
TABLE III: Average Waiting Time (s) at four unsignalized intersections under different traffic control scenarios. Increasing RV rates lead to significant reductions in waiting times across all intersections compared to HV-Signalized and HV-Unsignalized, especially when RV penetration rate is 60% or above. For intersection I, waiting time reduces from 122.35s at 10% RV rate to 12.90s at 90% RV rate. Intersection II achieves its best performance at 80% RV rate, with an 86.39% reduction compared to HV-Signalized. Intersections III and IV demonstrate more variable improvements, but still achieve significant reductions in waiting times at higher RV rates (e.g., 60% and 50%, respectively).
Intersection HV-Signalized HV-Unsignalized RV Penetration Rate
10% 20% 30% 40% 50% 60% 70% 80% 90%
I 37.37 127.97 122.35 97.16 90.85 47.78 38.52 21.70 23.31 17.31 12.90
II 61.13 25.54 31.88 38.13 11.72 16.77 9.54 8.34 9.63 8.32 8.81
III 50.61 112.94 105.07 81.71 62.95 46.51 50.15 25.24 39.94 44.46 31.05
IV 39.71 101.32 54.08 56.33 50.12 14.59 8.92 22.11 17.87 11.83 10.39
Table III highlights the results for average waiting time at four intersections under different scenarios. Across all four intersections, increasing the RV rate leads to substantial reductions in average waiting times compared to both baseline scenarios, indicating improved traffic efficiency. At intersection I, the waiting time decreases from 122.35 s at 10% RV to 12.90 s at 90% RV, achieving maximum improvements of 65.48% and 89.92% over HV-Signalized and HV-Unsignalized. At intersection II, the highest improvement comes at RV rate 80% (8.32 s), providing a 86.39% gain over HV-Signalized and 67.42% over HV-Unsignalized baselines. Intersections III and IV also demonstrate significant gains, with maximum improvements of 50.13% and 77.54% over signalized, and 77.65% and 91.19% over unsignalized baselines, respectively. These results highlight the potential of RVs in dramatically reducing congestion and improving throughput of mixed traffic, compared to traditional control method via traffic lights. Fig. 2 shows an example scenario at intersection I.
Figure 2: Traffic conditions at intersection I at 10-minute mark, RVs in red and HVs in white. (a) HV-Signalized: Traffic flow is controlled via traffic lights. (b) HV-Unsignalized: Intersection quickly becomes congested without traffic lights, resulting in severe gridlock in all directions. (c) Mixed traffic at 10% RV rate: While slightly better than HV-Unsignalized, HVs still dominate the flow, limiting the RVs’ ability to fully optimize the intersection. (d) Mixed traffic at 60% RV rate: RVs learn to adapt to unsignalized conditions, significantly reducing congestion and improving traffic flow. This highlights the potential of RV-based traffic management in complex intersections.
IV-C Traffic Heterogeneity Evaluation
The incorporation of RVs into traffic systems poses unique challenges when considering the diverse range of vehicle types present on urban roadways. These challenges arise from varying vehicle sizes, speeds, and maneuvering capabilities, which complicate traffic flow coordination including intersection management. Here we examine how different vehicle types—passenger cars, pickups, vans, semi-trailers, and trucks—respond to increasing RV rates in terms of traffic efficiency and environmental impact.
IV-C1 Average Waiting Time
Fig. 3 (top row) presents the average waiting times at intersections for different vehicle types across RV rates, compared to HV-Signalized. We observe that all vehicle types experience a decrease in waiting time as RV rate increases, indicating a broad positive impact of RV introduction on traffic efficiency. However, the magnitude of improvement varies across vehicle types.
Figure 3: Average Waiting Time (top row), CO2 Emissions (middle row), and Fuel Consumption (bottom row) for different vehicle types across RV rates (10% to 90%) compared to HV-Signalized. As RV rate increases, waiting times generally decrease for all vehicle types, outperforming HV-Signalized starting from 40% RV penetration. CO2 emissions and fuel consumption show slight increases but remain below HV-Signalized levels for all RV scenarios. Larger vehicles like trucks and semi-trailers show more dramatic increases in emissions and fuel consumption as RV rates rise, due to their size and weight. In contrast, passenger cars and other smaller vehicles exhibit moderate increases in environmental impact while still benefiting from significant waiting time reductions, demonstrating a balanced improvement across all metrics.
Trucks exhibit the most significant improvement, with waiting time reducing from 60.80 s at 10% RV rate to 4.50 s at 90% RV rate—an 87.03% reduction compared to HV-Signalized (34.68 s). Pickups, with the second lowest percentage in the traffic, show similar patterns of improvement—up to 77.52% reduction vs. HV-Signalized. Semi-trailers show less dramatic improvements, with waiting time reducing from 79.97 s to 20.12 s—61.57% reduction compared to HV-Signalized. Passenger cars and vans constitute the majority of traffic, showing significant but comparatively smaller improvements. Passenger cars see their average waiting time decrease to 23.99 s, while vans improve to 19.68 s, achieving 52.46% and 61.51% improvements, respectively, compared to HV-Signalized. The results show varied impacts of RV coordination, with less frequent vehicles like trucks and pickups experiencing the greatest reductions in waiting times, while more common vehicles like passenger cars and vans show moderate improvements.
TABLE IV: Percentage improvements in waiting times for vehicle types at different RV rates compared to HV-Signalized. Bold values indicate the highest improvement for each vehicle type, all occurring at 90% RV rate. This trend demonstrates rarity advantage for less common vehicles and the benefit of a more homogeneous environment for passenger cars. Trucks show the most significant improvement (87.03%), followed by pickups (77.52%). Passenger cars (both RV and non-RV), despite being the most common, still achieve substantial improvements (52.46%) due to the reduced heterogeneity of the traffic environment at high RV rates.
Vehicle Type RV Penetration Rate
40% 50% 60% 70% 80% 90%
Passenger Car 16.49 29.40 45.00 44.89 46.98 52.46
Pickup 41.40 50.36 69.09 57.46 60.67 77.52
Semi-trailer 33.55 39.16 54.63 50.95 58.41 61.57
Truck 60.22 62.14 80.30 56.88 65.02 87.03
Van 24.58 41.78 53.58 51.43 54.82 61.51
To understand this trend better, we present the percentage improvements in waiting times for the vehicle types at key RV rates compared to HV-Signalized in Table IV. The disproportionate benefit for less-frequent vehicles like trucks and pickups can be attributed to their ability to take advantage of the gaps created by RV coordination without contributing significantly to overall congestion. Their infrequent presence at intersections means they are more likely to encounter optimized traffic conditions created by the RVs. Similarly, semi-trailers’ moderate improvement compared to trucks might be due to their higher frequency in traffic, leading to more vehicle interactions and potentially more complex maneuvering requirements. In contrast, more common vehicles like passenger cars and vans see moderate improvements, likely due to their frequent interactions with other vehicles.
IV-C2 CO2 Emissions and Fuel Consumption
We present the trends in CO2 emissions and fuel consumption across different vehicle types as RV rate increases in Fig. 3 (middle and bottom rows).
For CO2 emissions, we observe a general increasing trend for all vehicle types as RV rates rise. Passenger cars show a moderate increase, with emissions rising from 2,333 mg/s at 10% RV rate to 2,917 mg/s at 90% RV rate, a 25% increase. Pickups exhibit a similar trend, experiencing a 24% rise, with emissions increasing from 2,261 mg/s to 2,813 mg/s. Larger vehicles demonstrate more substantial increases in CO2 emissions. Trucks, for example, see their emissions rise from 8,497 mg/s at 10% RV rate to 12,946 mg/s at 90% RV rate, leading to a 52% increase. Semi-trailers show the most dramatic change, with emissions going from 7,554 mg/s to 12,165 mg/s, a significant 61% increase. The overall trend shows that CO2 emissions rise more sharply for larger vehicles as RV rate increases.
Fuel consumption patterns closely mirror the CO2 emission trends. Semi-trailers show the most significant increase, with fuel consumption rising from about 3.2 ml/s at 10% RV rate to 5.2 ml/s at 90% RV rate, a 62.5% increase. Trucks follow a similar pattern, with consumption increasing from 3.6 ml/s to 5.6 ml/s, a 55.5% rise. Passenger cars, pickups, and vans exhibit lower overall fuel consumption but still show an increasing trend with higher RV rates. For instance, passenger car fuel consumption rises from 1.0 ml/s to 1.25 ml/s, a 25% increase as RV rate goes from 10% to 90%. Overall, larger vehicles experience the greatest rise in fuel consumption at higher RV rates, while smaller vehicles see more moderate increases.
The rise in both CO2 emissions and fuel consumption for all vehicle types could be attributed to the shift from intermittent idling to more continuous motion. While vehicles spend less time stationary, they travel at higher speeds, potentially increasing overall fuel consumption and emissions. Due to their weight and lower fuel efficiency at higher speeds, larger vehicles like trucks and semi-trailers are more sensitive to this change, resulting in a net increases in emissions and fuel consumption. However, both CO2 emissions and fuel consumption remain significantly lower than HV-Signalized across all vehicle types and RV rates. This suggests that while there is a trade-off between reduced waiting times and increased emission/fuel consumption as RV rate rises, the net impact of RV-managed intersections remains more environment-friendly than signalized intersections.
IV-C3 Space Headway
As shown in Fig. 4, we observe a consistent reduction in space headways across all vehicle types as RV rate increases. The impact of RVs on space headways varies across vehicle types. Passenger cars, which dominate the traffic composition, show a marked decrease in space headways from 57.49 m at 10% RV rate to 35.78 m at 90% RV rate, a reduction of about 38%. Larger vehicles such as semi-trailers, which typically require more space due to their size and slower acceleration, also exhibit reduced space headways, with space headways decreasing from 70.55 m to 53.21 m (25% reduction) as RV rate increases from 10% to 90%. These results underscore the potential of RVs in facilitating streamlined traffic flows.
Figure 4: Space headways for different vehicle types at varying RV penetration rates. As RV rate increases, space headways decrease across all vehicle types, leading to better road space utilization. Passenger cars show the largest reduction. Trucks and semi-trailers see smaller reductions, with a plateau effect at higher RV rates.
In case of less occurring vehicles (e.g., trucks and pickups) at higher RV rates, we observe a plateau effect, where headways stabilize or slightly increase. This could be attributed to the fact that low-frequency vehicles interact less frequently at intersections and face fewer congestion issues in high-RV scenarios. Therefore, they do not experience the same pressure to reduce headways as passenger cars do. The changing vehicle type mix at higher RV rates may also contribute to this effect.
V CONCLUSION AND FUTURE WORK
This study investigates the impact of heterogeneous mixed traffic control using RVs in complex environments. Our results show that higher RV rates significantly reduce average waiting times for all vehicle types, with reductions of up to 86% and 91% compared to HVs at signalized and unsignalized intersections, respectively. Notably, less common vehicles like trucks benefit the most from RV coordination. While RVs lead to increased emissions and fuel consumption due to higher speeds, the overall efficiency gains surpass those of signalized intersections. Additionally, we observe consistent reductions in space headways across all vehicle types, with the largest decrease for passenger cars.
There exist several future directions. First, we would like to expand the testing of heterogeneous mixed traffic to larger-scale networks by leveraging existing traffic simulation, reconstruction, and prediction techniques [40, 41, 42, 43, 44, 45, 46, 47]. Second, we plan to enhance the capabilities of RVs by integrating advanced autonomous features in both 2D simulation [48, 49, 50] and 3D simulation [51, 52]. Finally, testing communication strategies among RVs under both short and long communication ranges through network optimization [53] could further improve the applicability of our approach.